Shift Summary for owl shift, 29/06/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Tue, 29 Jun 1999 07:23:45 -0500

Shift Summary for owl shift, 29/06/99

On Shift: Elliott Cheu, Ashkan Alavi

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of E832 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Accelerator down, stop and start the runs,
ET problem, several Etotal map reloading.
========================================================================
Runs ended this shift:

Run 13756 Beam_832: 13.9M Events

Tapes brought to FCC this shift:
(including tapes for online split)
QKC055-57, QKK055-57, QKS055-57, QKB019

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

Relatively smooth BEAM_832 data taking. Problem with Etotal in the last
two hours. Occasional large spikes in the DC currents (up to 250 uA).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13756 35 0:00 No beam, investigating 8GeV line corrector problem.
54 0:28 Beam returns, Sem=8.4E12.
73 0:50 ktevana error (crate 53).
96 1:25 240uA spike in chamber 1.
116 1:50 laser crate DYC is full, KFREND.
3:45 More spikes in DCs occurring at the beginning of
spill (see logbook p85). Called MCR.
258 5:00 Elliott noticed CsI histograms were not updating.
260 5:05 Stop the run. It turned out that level3 filter
process had crashed at 4:18!

13757 5:15 On initializing this run, pipe06 does not respond.
Rebooted the pipeline.
4 5:25 Start the run.
10 5:40 ET shows stripes in TE quadrant. Reload ET.
Elliott starts a race to beat his voice in catching
the problem!
21 5:50 ET shows stripes in TE quadrant. Reload ET.
one process in ktev1 is taking a factor or 2 more
time!
28 6:00 ET shows stripes in TE quadrant. Reload ET.
30 6:05 Crate 53 DYC full, trigger hangs, KFREND.
32 6:08 ET shows stripes in TE quadrant. Reload ET.
33 6:10 ET shows stripes in TE quadrant. Reload ET.
49 6:29 KTEVANA error, KFREND.
55 6:36 ET shows stripes in TE quadrant. Reload ET.
62 6:47 ET shows stripes in TE quadrant. Reload ET.
66 6:51 ET shows stripes in TE quadrant. Reload ET.
68 6:56 ET shows stripes in TE quadrant. Reload ET.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Last evening shift experienced a problem with tape drive tps2d2. Jim
Graham has marked that as a bad drive and used tps3d2 for one of the
QKS tapes instead until the old one is fixed.
Keep an eye on the DC spikes.
The controller for Etot TE quadrant has to be replaced.